À magnet is parallel to a uniform magnetic field.if it is rotated by 53 o from the direction of magnetic field,the woek done is 16 J.how much work is done to move it 37o further from the direction of magnetic field?

Grade:12

1 Answers

Vikas TU
6 years ago
Work done = W = MB (cos θ1 – cos θ2)
            As per the question
ð 16 = MB (cos 0o – cos 53o)
ð MB =16/0.4 = 40
 
            Now the work done in moving 37o more would be
            W = 40 (cos 53o - cos 90 o) = 24 J